PLEASE HELP!!!
This is what the question says. I don't understand what they want me to do!
Write the first five terms of the sequence.
a_n = n^2 + 2

Answer by Aldorozos(172)

Replace n with numbers 1,2,3,4,5
Here is the first term
1^2+2= 1+2=3
Here is the second term
2^2 +2= 4+2= 6
Here is the third therm
3^2+2=11
And so on